# RefLoRA: Refactored Low-Rank Adaptation for Efficient Fine-Tuning of Large Models

### v1

@@ 1 Introduction
Large language models (LLMs) have revolutionized a wide spectrum of applications including chatbots [1], code generation [7], and scientific discovery [51]. Despite their success, adapting LLMs to specific tasks remains computationally demanding. LLMs are built on a two-stage learning process, namely pre-training and fine-tuning. Pre-training is performed on massive, Internet-scale corpora. This endows LLMs with in-text comprehension and generation abilities, but results in models with billions to trillions parameters [1; 16]. Though broad language capabilities are granted, pre-trained LLMs are yet not tailored for specialized applications. To acquire domain-specific expertise, LLMs must be further trained on downstream tasks, what is known as fine-tuning. With the continually growing model size however, conventional full fine-tuning approaches (optimizing all model parameters) can be increasingly prohibitive due to the immense GPU memory and substantial computational capacity demands, rendering them impossible for individual users and organizations.
To tackle the computational bottleneck, parameter-efficient fine-tuning (PEFT) [21] has been investigated to enhance fine-tuning efficiency. As opposed to fully fine-tuning all parameters, PEFT methods either optimize a sparse subset [17; 50], or introduce additional lightweight trainable parameters while keeping the pre-trained ones frozen [21; 44; 36; 28; 33]. Among these approaches, low-rank adaptation (LoRA) [22] has gained popularity due to its low additional cost during inference. LoRA presumes that the parameter updates during fine-tuning lie on a low-dimensional manifold, and thus can be captured by a low-rank matrix. Though effective, LoRA are observed to suffer from challenges such as slow and unstable convergence [41], inconsistent and unbalanced weight updates [63; 19], and notable performance gaps relative to full fine-tuning [22]. One key reason behind these challenges is the nonuniqueness of LoRA’s low-rank factorization [63].
@@
Related work. Building upon LoRA, several have been developed to ameliorate its performance. Some revise vanilla LoRA’s architecture or dynamically adjust its hyperparameters. For instance, DoRA [61] decomposes LoRA weights into magnitude and direction components to improve learning capacity and training stability. AdaLoRA [66] allocates per-layer rank on-the-fly to prune less salient updates. FedPara [24] (a.k.a. LoHa) and LoKr [62] respectively integrate the low-rank structure with Hadamard and Kronecker products for reduced communication cost and improved model expressiveness. Another line of work boosts LoRA’s behavior in early fine-tuning epochs through well-designed initialization. PiSSA [41] leverages truncated singular value decomposition (SVD) to keep the low-rank initialization close to the pre-trained weights, whereas LoRA-GA [56] aligns the first update direction with full fine-tuning. Other variants refine LoRA’s per-step optimization to promote empirical convergence. LoRA-Pro [57] redirects LoRA’s gradient to match the weight update of full fine-tuning. LoRA-RITE [63] substitutes the default Adam optimizer [25] with customized gradient calculation and moment estimation scheme. However, these gradient-altering strategies necessitate meticulous crafting to ensure stability and convergence. Our approach falls within the same family, but adheres strictly to standard backpropagation rule, avoiding direct gradient manipulation and requiring less computational overhead. To meet constraints of space limits, additional LoRA variants are discussed in Appendix A.
@@ 2 Preliminaries and nonunique low-rank factorizations
Consider a general weight matrix $\mathbf{W}\in^{m\times n}$ parameterizing a large model. With initialization being the pre-trained matrix $\mathbf{W}_{0}:=\mathbf{W}^{\mathrm{pre}}$ and $t$ indexing iteration, conventional full fine-tuning updates the sizable matrix $\mathbf{W}_{t}$ by backpropagating the loss function $\ell(\mathbf{W}_{t})$. Though straightforward, this approach requires an excessive memory footprint and major computational cost.
Aiming at efficient fine-tuning, LoRA [22] freezes $\mathbf{W}^{\mathrm{pre}}$ and presumes that the weight increment exhibits a low-rank structure $\mathbf{W}_{t}=\mathbf{W}^{\mathrm{pre}}+\mathbf{A}_{t}\mathbf{B}_{t}^{\top}$, where $\mathbf{A}_{t}\in^{m\times r}$, $\mathbf{B}_{t}\in^{n\times r}$, and $r\ll\min\{m,n\}$ is a preselected constant. Consequently, LoRA’s objective function boils down to
$\min_{\mathbf{A},\mathbf{B}}\mathcal{L}(\mathbf{A},\mathbf{B}):=\ell(\mathbf{W}^{\mathrm{pre}}+\mathbf{A}\mathbf{B}^{\top}).$
This reformulation reduces the number of trainable parameters to $(m+n)r\ll mn$, thus effectively minimizing the associated memory and computation costs. As for initialization, LoRA draws entries of $\mathbf{A}_{0}$ from a zero-mean normal distribution with small variance $\sigma^{2}$ for numerical stability; and $\mathbf{B}_{0}=\mathbf{0}$ to ensure $\mathbf{W}_{0}=\mathbf{W}^{\mathrm{pre}}$. This choice gives rise to imbalanced updates of $\mathbf{A}_{t}$ and $\mathbf{B}_{t}$, and decelerates empirical convergence especially in early epochs. At the first iteration for example, the chain rule implies that $\nabla_{\mathbf{A}_{0}}\mathcal{L}(\mathbf{A}_{0},\mathbf{B}_{0})=\nabla\ell(\mathbf{W}^{\mathrm{pre}}+\mathbf{A}_{0}\mathbf{B}_{0}^{\top})\mathbf{B}_{0}=\mathbf{0}$, meaning no update to $\mathbf{A}_{0}$. In comparison, $\mathbf{B}_{0}$’s gradient $\nabla_{\mathbf{B}_{0}}\mathcal{L}(\mathbf{A}_{0},\mathbf{B}_{0})=\nabla\ell(\mathbf{W}^{\mathrm{pre}})^{\top}\mathbf{A}_{0}$ is generally non-zero. In fact, it turns out that $\|\mathbf{B}_{t}\|_{\mathrm{F}}\ll\|\mathbf{A}_{t}\|_{\mathrm{F}}$ and $\|\nabla_{\mathbf{A}_{t}}\mathcal{L}(\mathbf{A}_{t},\mathbf{B}_{t})\|_{\mathrm{F}}\ll\|\nabla_{\mathbf{B}_{t}}\mathcal{L}(\mathbf{A}_{t},\mathbf{B}_{t})\|_{\mathrm{F}}$ when $t$ is small [41], and these unbalanced low-rank factors and updates can markedly impede the convergence of LoRA.
@@
Notation.  Bold lowercase (capital) letters denote vectors (matrices); $\|\cdot\|_{2}$, $\|\cdot\|_{*}$, $\|\cdot\|_{\mathrm{F}}$ and $\langle\cdot,\cdot\rangle_{\mathrm{F}}$ stand for $\ell_{2}$-, nuclear-, Frobenius-norm and Frobenius inner product; $\operatorname{Col}(\cdot)$, $\operatorname{Null}(\cdot)$ and $[\cdot]_{i}$ represent column space, null space and the $i$-th entry/column of a vector/matrix; $\cdot^{\dagger}$, $\operatorname{tr}(\cdot)$, and $\operatorname{rank}(\cdot)$ refer to the Moore-penrose pseudoinverse, trace, and rank; $\lambda_{i}(\cdot)$ and $\sigma_{i}(\cdot)$ are the $i$-th largest eigenvalue and singular value; $\mathbb{S}_{++}^{r}$ indicates the set of $r\times r$ symmetric positive definite (SPD) matrices; $\mathrm{GL}(r)$ denotes the general linear group of degree $r$ (i.e., the set of $r\times r$ invertible matrices); and $\mathrm{O}(r)$ stands for the orthogonal group of degree $r$ (i.e., the set of $r\times r$ orthogonal matrices).
@@ 3.1 Characterizing LoRA’s factorization and weight update
Our analysis begins with the following mild assumption, which has been utilized and validated on various realistic datasets [57].
@@
Moreover, if $\>\mathbf{P}_{t}\in\mathrm{O}(r)$, then $\Delta\mathbf{W}_{t}=\Delta\tilde{\mathbf{W}}_{t}$.
@@ 3.2 Optimizing $\mathbf{S}_{t}$ via loss upper bound minimization
The loss function $\ell$ has $L$-Lipschitz gradient; i.e., $\|\nabla\ell(\mathbf{W})-\nabla\ell(\mathbf{W}^{\prime})\|_{\mathrm{F}}\leq L\|\mathbf{W}-\mathbf{W}^{\prime}\|_{\mathrm{F}},\;\forall\mathbf{W},\mathbf{W}^{\prime}\in^{m\times n}$.
Assumption 2 is equivalent to requiring Lipschitz smoothness of $\ell$ w.r.t. the vectorized weight matrix $\textrm{vec}(\mathbf{W})$, which is fairly mild and common in machine learning [48; 15] and optimization [3; 5]. Under this assumption, the loss admits the following quadratic upper bound
@@
Although $\ell$ is often Lipschitz smooth w.r.t. $\mathbf{W}_{t}$, its smoothness constants w.r.t. $\mathbf{A}_{t}$ and $\mathbf{B}_{t}$ can be unbounded due to the bilinear structure, unless one assumes boundedness or convergence of $\mathbf{A}_{t}$ and $\mathbf{B}_{t}$ [12]. Since these two assumptions are overly restrictive, they will be avoided in our analysis.
The nonlinear dependence of $\Delta\tilde{\mathbf{W}}_{t}$ on $\mathbf{S}_{t}$ (cf. (5)) prevents an analytical solution when directly optimizing $\mathbf{S}_{t}$ over the quadratic upper bound of $\ell(\mathbf{W}_{t}+\Delta\tilde{\mathbf{W}}_{t}(\mathbf{S}_{t}))$. This motivates iterative solvers involving matrix multiplication with the sizable matrix $\nabla\ell(\mathbf{W}_{t})\in^{m\times n}$. To mitigate the overhead, the next proposition relaxes the quadratic upper bound to factor out $\|\nabla\ell(\mathbf{W}_{t})\|_{2}^{2}$, thus decoupling $\mathbf{S}_{t}$’s optimization from $\nabla\ell(\mathbf{W}_{t})$.
@@
The high-order term $\mathcal{O}(L\eta^{3})$ originates from $\Delta\mathbf{A}_{t}\Delta\mathbf{B}_{t}^{\top}$ in (5). As $\eta$ is typically small ($\sim\mathcal{O}(10^{-4})$), this term is negligible in practice [56; 63]. As a consequence, the upper bound in (7) is dominated by its first term. This leads to our RefLoRA objective
@@
Theorem 3 states that if $\eta>0$ is not too small, then (8) admits a unique global optimum $\tilde{\mathbf{S}}_{t}$. Otherwise, there can be multiple optima, while two can always be constructed by appropriately scaling $\tilde{\mathbf{S}}_{t}$. In addition, $\eta<0$ is included in (10) for visualization purposes. We remark that $\tilde{\mathbf{S}}_{t}$ in (9) is known as the matrix geometric mean $(\mathbf{A}_{t}^{\top}\mathbf{A}_{t})^{-1}\#(\mathbf{B}_{t}^{\top}\mathbf{B}_{t})$ of $(\mathbf{A}_{t}^{\top}\mathbf{A}_{t})^{-1}$ and $(\mathbf{B}_{t}^{\top}\mathbf{B}_{t})$ [32], which can be also written as $\tilde{\mathbf{S}}_{t}=(\mathbf{A}_{t}^{\top}\mathbf{A}_{t})^{-1}[\mathbf{A}_{t}^{\top}\mathbf{A}_{t}\mathbf{B}_{t}^{\top}\mathbf{B}_{t}]^{\frac{1}{2}}$; cf. Lemma 7 in the Appendix.
@@
Figure 1: Visualization of loss $\ell(\mathbf{W}_{t}+\Delta\tilde{\mathbf{W}}_{t})$, quadratic (quad.) bound (6), and alternative (alt.) bound (7). LoRA corresponds to $\mathbf{S}_{t}=\mathbf{I}_{r}$, while our refactoring (ref.) optimizes $\mathbf{S}_{t}$.
Figure 1 plots the loss $\ell(\mathbf{W}_{t}+\Delta\tilde{\mathbf{W}}_{t})$, the quadratic upper bound (6) (with $\Delta\mathbf{W}_{t}$ replaced by $\Delta\tilde{\mathbf{W}}_{t}$), and the alternative upper bound (7) of a numerical example as a function of $\eta$; see also Appendix D.2 for details. LoRA corresponds to the non-optimized $\mathbf{S}_{t}=\mathbf{I}_{r}$, whereas our refactoring selects $\mathbf{S}_{t}=\mathbf{S}_{t}^{*}$ based on Theorem 3. Notably, $\eta=0$ leads to a jump discontinuity for our refactoring. Figure 1(a) shows that by optimizing the alternative bound, the associated quadratic bound and loss both become lower and flatter. This enables a larger step size $\eta$ to achieve a lower loss, thereby accelerating the empirical convergence of LoRA; see Section 4 for experiments corroborating this claim.
The flatter loss landscape arises thanks to balancing $\tilde{\mathbf{A}}_{t}$ and $\tilde{\mathbf{B}}_{t}$. Indeed, the imbalance of $\mathbf{A}_{t}$ and $\mathbf{B}_{t}$ in vanilla LoRA necessitates different step sizes [19]. Appendix B.3 proves that the balance $\tilde{\mathbf{A}}_{t}^{\top}\tilde{\mathbf{A}}_{t}=\tilde{\mathbf{B}}_{t}^{\top}\tilde{\mathbf{B}}_{t}$ is guaranteed with $\mathbf{S}_{t}=\tilde{\mathbf{S}}_{t}$, thus enabling a unified step size. With $\eta$ too small however, it is more beneficial (compared to balanced updates) to scale either $\mathbf{A}_{t}$ or $\mathbf{B}_{t}$ to accommodate the small $\eta$. This corresponds to the two solutions in the $0<\eta<1/(\tilde{C}_{t}L)$ case of (10).

### v2

@@ 1 Introduction
Large language models (LLMs) have revolutionized a wide spectrum of applications including chatbots GPT4 , code generation Codex , and scientific discovery Galactica . Despite their success, adapting LLMs to specific tasks remains computationally demanding. LLMs are built on a two-stage learning process, namely pre-training and fine-tuning. Pre-training is performed on massive, Internet-scale corpora. This endows LLMs with in-text comprehension and generation abilities, but results in models with billions to trillions parameters GPT4 ; Llama3 . Though broad language capabilities are granted, pre-trained LLMs are yet not tailored for specialized applications. To acquire domain-specific expertise, LLMs must be further trained on downstream tasks, what is known as fine-tuning. With the continually growing model size however, conventional full fine-tuning approaches (optimizing all model parameters) can be increasingly prohibitive due to the immense GPU memory and substantial computational capacity demands, rendering them impossible for individual users and organizations.
To tackle the computational bottleneck, parameter-efficient fine-tuning (PEFT) adapter-H has been investigated to enhance fine-tuning efficiency. As opposed to fully fine-tuning all parameters, PEFT methods either optimize a sparse subset diff-pruning ; sparse-mask , or introduce additional lightweight trainable parameters while keeping the pre-trained ones frozen adapter-H ; adapter-D ; adapter-L ; prompt-tuning ; prefix-tuning . Among these approaches, low-rank adaptation (LoRA) LoRA has gained popularity due to its low additional cost during inference. LoRA presumes that the parameter updates during fine-tuning lie on a low-dimensional manifold, and thus can be captured by a low-rank matrix. Though effective, LoRA is observed to suffer from challenges such as slow and unstable convergence PiSSA , inconsistent and unbalanced weight updates LoRA-RITE ; LoRA+ , and notable performance gaps relative to full fine-tuning LoRA . One key reason behind these challenges is the nonuniqueness of LoRA’s low-rank factorization LoRA-RITE .
@@
Related work. Building upon LoRA, several have been developed to ameliorate its performance. Some revise vanilla LoRA’s architecture or dynamically adjust its hyperparameters. For instance, DoRA DoRA decomposes LoRA weights into magnitude and direction components to improve learning capacity and training stability. AdaLoRA AdaLoRA and GeoLoRA GeoLoRA allocates per-layer rank on-the-fly to prune less salient updates. FedPara LoHa (a.k.a. LoHa) and LoKr LoKr respectively integrate the low-rank structure with Hadamard and Kronecker products for reduced communication cost and improved model expressiveness. Another line of work boosts LoRA’s behavior in early fine-tuning epochs through well-designed initialization. PiSSA PiSSA leverages truncated singular value decomposition (SVD) to keep the low-rank initialization close to the pre-trained weights, whereas LoRA-GA LoRA-GA aligns the first update direction with full fine-tuning. Other variants refine LoRA’s per-step optimization to promote empirical convergence. LoRA-Pro LoRA-Pro redirects LoRA’s gradient to match the weight update of full fine-tuning. LoRA-RITE LoRA-RITE substitutes the default Adam optimizer Adam with customized gradient calculation and moment estimation scheme. However, these gradient-altering strategies necessitate meticulous crafting to ensure stability and convergence. Our approach falls within the same family, but adheres strictly to standard backpropagation rule, avoiding direct gradient manipulation and requiring less computational overhead. To meet constraints of space limits, additional LoRA variants are discussed in Appendix A.
@@ 2 Preliminaries and nonunique low-rank factorizations
Consider a general weight matrix $\mathbf{W}\in\mathbb{R}^{m\times n}$ parameterizing a large model. With initialization being the pre-trained matrix $\mathbf{W}_{0}:=\mathbf{W}^{\mathrm{pt}}$ and $t$ indexing iteration, conventional full fine-tuning updates the sizable matrix $\mathbf{W}_{t}$ by backpropagating the loss function $\ell(\mathbf{W}_{t})$. Though straightforward, this approach requires an excessive memory footprint and major computational cost.
Aiming at efficient fine-tuning, LoRA LoRA freezes $\mathbf{W}^{\mathrm{pt}}$ and presumes that the weight increment exhibits a low-rank structure $\mathbf{W}_{t}=\mathbf{W}^{\mathrm{pt}}+\mathbf{A}_{t}\mathbf{B}_{t}^{\top}$, where $\mathbf{A}_{t}\in\mathbb{R}^{m\times r}$, $\mathbf{B}_{t}\in\mathbb{R}^{n\times r}$, and $r\ll\min\{m,n\}$ is a preselected constant. Consequently, LoRA’s objective function boils down to
$\min_{\mathbf{A},\mathbf{B}}\mathcal{L}(\mathbf{A},\mathbf{B}):=\ell(\mathbf{W}^{\mathrm{pt}}+\mathbf{A}\mathbf{B}^{\top}).$
This reformulation reduces the number of trainable parameters to $(m+n)r\ll mn$, thus effectively minimizing the associated memory and computation costs. As for initialization, LoRA draws entries of $\mathbf{A}_{0}$ from a zero-mean normal distribution with small variance $\sigma^{2}$ for numerical stability; and $\mathbf{B}_{0}=\mathbf{0}$ to ensure $\mathbf{W}_{0}=\mathbf{W}^{\mathrm{pt}}$. This choice gives rise to imbalanced updates of $\mathbf{A}_{t}$ and $\mathbf{B}_{t}$, and decelerates empirical convergence especially in early epochs. At the first iteration for example, the chain rule implies that $\nabla_{\mathbf{A}_{0}}\mathcal{L}(\mathbf{A}_{0},\mathbf{B}_{0})=\nabla\ell(\mathbf{W}^{\mathrm{pt}}+\mathbf{A}_{0}\mathbf{B}_{0}^{\top})\mathbf{B}_{0}=\mathbf{0}$, meaning no update to $\mathbf{A}_{0}$. In comparison, $\mathbf{B}_{0}$’s gradient $\nabla_{\mathbf{B}_{0}}\mathcal{L}(\mathbf{A}_{0},\mathbf{B}_{0})=\nabla\ell(\mathbf{W}^{\mathrm{pt}})^{\top}\mathbf{A}_{0}$ is generally non-zero. In fact, it turns out that $\|\mathbf{B}_{t}\|_{\mathrm{F}}\ll\|\mathbf{A}_{t}\|_{\mathrm{F}}$ and $\|\nabla_{\mathbf{A}_{t}}\mathcal{L}(\mathbf{A}_{t},\mathbf{B}_{t})\|_{\mathrm{F}}\ll\|\nabla_{\mathbf{B}_{t}}\mathcal{L}(\mathbf{A}_{t},\mathbf{B}_{t})\|_{\mathrm{F}}$ when $t$ is small PiSSA , and these unbalanced low-rank factors and updates can markedly impede the convergence of LoRA.
@@
Notation.  Bold lowercase (capital) letters denote vectors (matrices); $\|\cdot\|_{2}$, $\|\cdot\|_{*}$, $\|\cdot\|_{\mathrm{F}}$ and $\langle\cdot,\cdot\rangle_{\mathrm{F}}$ stand for $\ell_{2}$-, nuclear-, Frobenius-norm and Frobenius inner product; $\operatorname{Col}(\cdot)$, $\operatorname{Null}(\cdot)$ and $\cdot]_{i}$ represent column space, null space and the $i$-th entry/column of a vector/matrix; $\cdot^{\dagger}$, $\operatorname{tr}(\cdot)$, and $\operatorname{rank}(\cdot)$ refer to the Moore-penrose pseudoinverse, trace, and rank; $\lambda_{i}(\cdot)$ and $\sigma_{i}(\cdot)$ are the $i$-th largest eigenvalue and singular value; $\mathbb{S}_{++}^{r}$ indicates the set of $r\times r$ symmetric positive definite (SPD) matrices; $\mathrm{GL}(r)$ denotes the general linear group of degree $r$ (i.e., the set of $r\times r$ invertible matrices); and $\mathrm{O}(r)$ stands for the orthogonal group of degree $r$ (i.e., the set of $r\times r$ orthogonal matrices).
@@ 3.1 Characterizing LoRA’s factorization and weight update
Our analysis begins with the following mild assumption, which has been utilized and validated on various realistic datasets LoRA-Pro .
@@
Moreover, if $\mathbf{P}_{t}\in\mathrm{O}(r)$, then $\Delta\tilde{\mathbf{W}}_{t}=\Delta\mathbf{W}_{t}$.
@@ 3.2 Optimizing $\mathbf{S}_{t}$ via loss upper bound minimization
The loss function $\ell$ has $L$-Lipschitz gradient; i.e., $\|\nabla\ell(\mathbf{W})-\nabla\ell(\mathbf{W}^{\prime})\|_{\mathrm{F}}\leq L\|\mathbf{W}-\mathbf{W}^{\prime}\|_{\mathrm{F}},\;\forall\mathbf{W},\mathbf{W}^{\prime}\in\mathbb{R}^{m\times n}$.
Assumption 2 is equivalent to requiring Lipschitz smoothness of $\ell$ w.r.t. the vectorized weight matrix $\textrm{vec}(\mathbf{W})$, which is fairly mild and common in machine learning Understand-ML ; DL-Goodfellow-Bengio and optimization NP-Bertsekas ; cvx-opt-Boyd . Under this assumption, the loss admits the following quadratic upper bound
@@
Although $\ell$ is often Lipschitz smooth w.r.t. $\mathbf{W}_{t}$, its smoothness constants w.r.t. $\mathbf{A}_{t}$ and $\mathbf{B}_{t}$ can be unbounded due to the bilinear structure, unless one assumes boundedness or convergence of $\mathbf{A}_{t}$ and $\mathbf{B}_{t}$ MF-GD . Since these two assumptions are overly restrictive, they will be avoided in our analysis.
The nonlinear dependence of $\Delta\tilde{\mathbf{W}}_{t}$ on $\mathbf{S}_{t}$ (cf. (5)) prevents an analytical solution when directly optimizing $\mathbf{S}_{t}$ over the quadratic upper bound of $\ell(\mathbf{W}_{t}+\Delta\tilde{\mathbf{W}}_{t}(\mathbf{S}_{t}))$. This motivates iterative solvers involving matrix multiplication with the sizable matrix $\nabla\ell(\mathbf{W}_{t})\in\mathbb{R}^{m\times n}$. To mitigate the overhead, the next proposition relaxes the quadratic upper bound to factor out $\|\nabla\ell(\mathbf{W}_{t})\|_{2}^{2}$, thus decoupling $\mathbf{S}_{t}$’s optimization from $\nabla\ell(\mathbf{W}_{t})$.
@@
The high-order term $\mathcal{O}(L\eta^{3})$ originates from $\Delta\mathbf{A}_{t}\Delta\mathbf{B}_{t}^{\top}$ in (5). As $\eta$ is typically small ($\sim\mathcal{O}(10^{-4})$), this term is negligible in practice LoRA-GA ; LoRA-RITE . As a consequence, the upper bound in (7) is dominated by its first term. This leads to our RefLoRA objective
@@
Theorem 3 states that if $\eta>0$ is not too small, then (8) admits a unique global optimum $\tilde{\mathbf{S}}_{t}$. Otherwise, there can be multiple optima, while two can always be constructed by appropriately scaling $\tilde{\mathbf{S}}_{t}$. In addition, $\eta<0$ is included in (10) for visualization purposes. We remark that $\tilde{\mathbf{S}}_{t}$ in (9) is known as the matrix geometric mean $(\mathbf{A}_{t}^{\top}\mathbf{A}_{t})^{-1}\#(\mathbf{B}_{t}^{\top}\mathbf{B}_{t})$ of $(\mathbf{A}_{t}^{\top}\mathbf{A}_{t})^{-1}$ and $(\mathbf{B}_{t}^{\top}\mathbf{B}_{t})$ geo-mean , which can be also written as $\tilde{\mathbf{S}}_{t}=(\mathbf{A}_{t}^{\top}\mathbf{A}_{t})^{-1}[\mathbf{A}_{t}^{\top}\mathbf{A}_{t}\mathbf{B}_{t}^{\top}\mathbf{B}_{t}]^{\frac{1}{2}}$; cf. Lemma 7 in the Appendix.
@@
Figure 1: Visualization of loss $\ell(\mathbf{W}_{t}+\Delta\tilde{\mathbf{W}}_{t})$ and upper bound (7). LoRA corresponds to $\mathbf{S}_{t}=\mathbf{I}_{r}$, while our refactoring (ref.) optimizes $\mathbf{S}_{t}$.
Figure 1 plots the loss $\ell(\mathbf{W}_{t}+\Delta\tilde{\mathbf{W}}_{t})$, and the upper bound (7) of a numerical example as a function of $\eta$; see also Appendix D.2 for details. LoRA corresponds to the non-optimized $\mathbf{S}_{t}=\mathbf{I}_{r}$, whereas our refactoring selects $\mathbf{S}_{t}=\mathbf{S}_{t}^{*}$ based on Theorem 3. Notably, $\eta=0$ leads to a jump discontinuity for our refactoring. Figure 1(a) shows that by optimizing the upper bound (7), the associated loss becomes lower and flatter. This enables a larger step size $\eta$ to achieve a lower loss, thereby accelerating the empirical convergence of LoRA; see Section 4 for experiments corroborating this claim.
The flatter loss landscape arises thanks to balancing $\tilde{\mathbf{A}}_{t}$ and $\tilde{\mathbf{B}}_{t}$. Indeed, the imbalance of $\mathbf{A}_{t}$ and $\mathbf{B}_{t}$ in vanilla LoRA necessitates different step sizes LoRA+ . Appendix B.3 proves that the balance $\tilde{\mathbf{A}}_{t}^{\top}\tilde{\mathbf{A}}_{t}=\tilde{\mathbf{B}}_{t}^{\top}\tilde{\mathbf{B}}_{t}$ is guaranteed with $\mathbf{S}_{t}=\tilde{\mathbf{S}}_{t}$, thus enabling a unified step size. With $\eta$ too small however, it is more beneficial (compared to balanced updates) to scale either $\mathbf{A}_{t}$ or $\mathbf{B}_{t}$ to accommodate the small $\eta$. This corresponds to the two solutions in the $0<\eta<1/(\tilde{C}_{t}L)$ case of (10).